when Mrs jones was older than 50 but younger than 80, she told a friend, "each of my sons has as many sons as brothers, and the number of my years is exactly that of the number of both my sons and grandsons."
Answer by ikleyn(52803)

.
when Mrs jones was older than 50 but younger than 80, she told a friend,
"each of my sons has as many sons as brothers, and the number of my years
is exactly that of the number of both my sons and grandsons."
~~~~~~~~~~~~~~~~~~

Let n be the number of her sons.


Each of the sons has (n-1) brothers.


So, to answer the problem's question, we should solve this inequality


    50 < n*(n-1) + n < 80


or, equivalently,


    50 < n^2 < 80,


which gives a unique solution  n = 8.


So, her age is  n*(n-1) + n = n^2 = 8^2 = 64.    ANSWER

Solved, answered, carefully explained and completed.